OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. y - 2 = 4(x + 6) y - 6 = 4(x + 2) y + 2 = 4(x - 6) y + 6 = 4(x - 2) I need someone to explain :(

OpenStudy (anonymous):

general equation line in point slope form is \[\Large y-y_{0}=m(x-x_{0})\] ok let me give you an example if a line passes through (-5,-2) with slope 5 then the line will be \[\Large y-(-2)=5(x-(-5))\] oR \[\Large \implies y+2=5(x+5)\] can you find it now !!!

OpenStudy (anonymous):

@pam142 can you find it now?

OpenStudy (anonymous):

yes thank you i am trying now

OpenStudy (anonymous):

ok i am here. when you find it out let me know your answer !

OpenStudy (anonymous):

I don't get this :'(

OpenStudy (anonymous):

so it's y-6=-2(x+4)

OpenStudy (anonymous):

no it's y-6=4(x-2) :D

OpenStudy (anonymous):

sorry !! you missed that slope is =4 you used -2 at the place of slope . y-6=4(x-(-2)) y-6=4(x+2)!

OpenStudy (anonymous):

OH so it's plus two because negative times a negative I get it

OpenStudy (anonymous):

so it is ok now !?

OpenStudy (anonymous):

Yeah thanks! uhm do you think you can help me with another one?

OpenStudy (anonymous):

yeah sure !

OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? x = (y + 3) y - 4x = -3 y = 4x - 3 -4x = -y - 3

OpenStudy (anonymous):

ok let me tel you general form of slope intercept form, it is \[\Large y=mx+C\] can you figure it out which one is in point slope form ?

OpenStudy (anonymous):

the third one?

OpenStudy (anonymous):

yes you are right ..well done.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you 're welcome:)